# Cognitive Memory in Large Language Models

### v1

@@ 3.2.4 Handling Contradictory Memories
In long - term conversations, the management and construction of memory is a complex and important task, especially when it comes to conflicting memories. It is commonly believed that avoiding conflicting memories is necessary Kim et al. (2024). However, this view does not entirely conform to the true characteristics of human cognition. In fact, human memory is highly context - dependent. The same person may recall different information in different contexts, and even memories of the same event may vary with changes in context. Therefore, allowing memories with conflicting interpretations to coexist within an individual’s memory system not only aligns with the natural laws of human cognition but also adds richness and authenticity to conversations. To effectively deal with these conflicting memories, the author has proposed several optimization strategies for large - language models to choose from flexibly in long - term conversations. First is the “Memory Resolution” strategy. Inspired by entity resolution methods, this strategy merges conflicting memories into an information - rich sentence based on the context of the memory sources, thereby cleverly resolving these conflicts. Second is the “Memory Disambiguation” strategy. The contradiction between two statements may arise from the absence of context, that is, pragmatic ambiguity. Drawing on entity disambiguation methods, the memory disambiguation strategy clarifies the true meaning of each memory by extracting relevant information from the context of the memory, thereby eliminating ambiguity. Finally, there is the “Retention” strategy. Due to the limitations of natural language inference (NLI) models, sometimes memories predicted to be contradictory may not actually be so, but rather a deviation in the model’s understanding. In such cases, retaining these seemingly conflicting memories can actually provide more possibilities and depth to the conversation. Through these strategies, the author aims to transform conflicting memories into sentences containing richer information about the speaker, thereby enhancing the quality and coherence of long - term conversations.
@@ 3.3.1 Full-text Search
Full-text search Chen et al. (2024b) involves scanning the entire text dataset to locate specific character sequences or strings. This method not only looks for exact matches but also considers approximate matches based on text structure and wording. In MAS, when facing broad queries or when the exact location or label of information is unknown, the agent can use full-text search. For example, when a user asks a broad question about “the impact of climate change,” the agent can utilize full-text search to browse its entire memory and extract relevant paragraphs or interactions.
@@ 3.3.5 Hash-based Search
Liu et al. (2023) proposed hash-based search methods. In the retrieval task based on memory storage, in order to obtain the most relevant thoughts, the author designed a two-stage retrieval paradigm. Specifically, when faced with a new query $Q$, the embedding vector $x$ of the query is first obtained using a large language model agent. Subsequently, the retrieval process based on Locality-Sensitive Hashing (LSH) comes into play. By applying the LSH function,
$\text{hash}(x)=\text{LSH}(x)$
@@ 3.3.6 Multiple Searches
In the memory retrieval process, the author employs a dense retrieval method, utilizing the SimCSE model to encode both user input and each memory record into vectors. By calculating the cosine similarity, the top K relevant records are selected for response generation. However, since the retrieved memory records may not all be useful and could even be completely irrelevant, the author introduces a self-reflection mechanism for memory. Specifically, after obtaining the retrieval results, the language model is first required to determine whether this information is sufficient to respond to the user’s input. If it is sufficient, the language model extracts the relevant parts from the memory records as the output. If not, the original query needs to be improved by generating keywords or phrases that represent the missing information, which are then expanded into the original query for another round of retrieval. This self-reflection process can be repeated multiple times until the language model finds sufficient information.
@@ 4 KV Cache-Based Memory
This chapter introduces long-term memory based on KV cache. Since transformer-based LLMs are currently the mainstream, especially the dominance of the decoder-only structure, KV cache-based memory has become a hot research topic, particularly in the field of long context. In terms of the structure of this chapter, we follow the structure of Chapter 3, introducing memory from three aspects: acquisition, management, and utilization. Since there are many research works related to memory acquisition, we divide it into two sections: KV selection and KV compression. KV selection refers to choosing some KVs to save and discarding the rest. We categorize the selection strategies into eight types: regularity-based summarization, score-based methods, special token embeddings, learning-based KV selection, different selection strategies for different layers and heads, identical KV cache for different layers and heads, locality-sensitive hashing strategy, and backtracking-based KV selection. For KV compression in memory acquisition, we summarize it as low-rank compression, KV merging, multimodal compression, KV quantization, and multi-level compression.
For the second major part, memory management, we divide it into offload strategies (that is, placing memory on cheaper memory), optimization strategies combined with the OS system, storage data structures, and shared attention. We also introduce an article that explores the relationship between KV cache compression and attention.
@@ 4.1.1 Strategies of Regularities
Liu et al. (2024b) speculate that not all tokens need to be stored in memory for large language models to understand context. Just as humans can skim an article and grasp its main ideas, large language models may also be able to skim and understand. It is commonly observed that the attention scores of a token follow a strong power-law distribution, meaning that a token only pays high attention to a few other tokens. More importantly, the authors have observed repeated attention patterns among different tokens in a sequence in trained large language models. Some tokens are more important throughout a paragraph. Specifically, for two different tokens, there is similarity in the objects they highly attend to, as well as in the objects they ignore. Inspired by these observations, the authors propose the hypothesis of importance persistence: Only those tokens that have had a significant impact in the previous step will have a notable impact in future steps. If this hypothesis holds true, it implies that it is possible to anticipate which tokens may be important for future generation.
@@ 4.1.3 Score Strategy
One important feature of the Quest algorithm Tang et al. (2024b) is its ability to dynamically adjust the contents of the KV cache. Specifically, as the query vector Q changes, Quest re-estimates the criticality of tokens and updates the contents of the KV cache accordingly. This dynamic adjustment ensures that the KV cache always contains the key information required for the current query. The authors’ insight is that to avoid missing critical tokens, they should select pages that contain tokens with the highest attention weights. However, to efficiently select pages, they should compute an approximate attention score based on this insight. The authors found that the upper bound of attention weights within a page can be used to approximate the highest attention within that page. The upper bound of attention weights can be calculated using the channel-wise minimum (mi) and maximum (Mi) of the key vectors. Given a query vector Q, Quest computes the maximum possible value for channel i by taking Ui = max(Qimi, QiMi). Note that Ui is always greater than the product of Qi and any key value Ki in that page, regardless of the sign of Qi. Therefore, when the authors sum Ui, they obtain the upper bound of attention weights for all key vectors in that page. After deriving the upper bound of attention weights, the authors select the top K pages as key pages, where K is an arbitrarily defined hyperparameter. To demonstrate the feasibility of Quest, the authors performed actual self-attention calculations and collected the Top-K attention scores for each page. As shown in Figure 3, the authors’ query-aware sparsity mostly aligns with the predicted sparsity. Quest only performs normal self-attention on the selected pages, which significantly reduces memory movement. The authors define the number of tokens in the selected pages as the "token budget."
@@
Sparse Window Attention (SWA) Zhao et al. (2024) is a method used in the inference process of large language models (LLMs), aiming to reduce memory usage by generating sparse attention patterns while maintaining model accuracy. The core of SWA lies in combining local static and global dynamic sparse patterns: it generates static patterns on local tokens, retaining the most recent tokens to maintain the sequential semantics of language, while capturing the dynamically changing semantic importance of previous tokens through dynamic patterns. This hybrid pattern can more effectively capture key tokens in the sequence.
@@
4.1.4 Special Token Embedding Strategies
@@
4.1.5 Learning-based Token Selection Strategies
@@
Wang et al. (2024b) mainly introduces an algorithm called SQUEEZEATTENTION, which is used to manage KV caches in LLM (Large Language Model) inference to further reduce the memory consumption of inference. The paper points out that existing KV cache optimization methods mostly focus on optimizing from the sequence dimension or the batch dimension, while neglecting the potential opportunities in the attention layer dimension. By analyzing the behavior of different attention layers during inference, the authors propose a method for precise re-allocation of cache budgets for attention layers, thereby further reducing the total amount of KV caches.
@@
Fu et al. (2024) propose Mixture of Attention (MoA), which automatically customizes different sparse attention configurations for various heads and layers. MoA constructs and navigates a search space of various attention patterns and their scaling rules relative to input sequence lengths. It profiles the model, evaluates potential configurations, and identifies the optimal sparse attention compression plan. MoA adapts to different input sizes, revealing that some attention heads expand their focus to accommodate longer sequences, while other heads consistently concentrate on fixed-length local contexts.
@@
SnapKV Li et al. (2024) is a KV cache compression method for optimizing large language models (LLMs) when processing long contexts. It is implemented through the following steps: First, an observation window located at the end of the prompt is used to monitor the model’s attention patterns on input tokens during the generation process, which are stable and consistent throughout generation. Next, a voting algorithm based on the observation window is employed to identify important positions that are consistently attended to by attention heads during generation, thereby selecting the KV pairs corresponding to the most important tokens. Before voting, SnapKV introduces a pooling-based clustering technique to capture complete local information and filter out irrelevant tokens, thus compressing the context while preserving its integrity. Finally, the selected KV pairs are concatenated with the observation window to form a new, smaller KV cache. The model only needs to perform attention calculations on this significantly reduced cache during subsequent generation, thereby improving efficiency.
@@
4.1.7 Different layers and different heads use the same token selection strategy.
@@ 4.2.1 Low-Rank Compression
KV-Cache compression methods typically sample a subset of effective tokens or quantize the data into lower numerical bit widths. However, these methods cannot exploit the redundancy in the hidden dimension of the KV tensors. Chang et al. (2024) presents a hidden dimension compression approach called Palu, a novel KV-Cache compression framework that leverages low-rank projection. Palu decomposes linear layers into low-rank matrices, caches the compressed intermediate states, and dynamically reconstructs the full keys and values. To improve accuracy, compression rate, and efficiency, Palu further includes a medium-grained low-rank decomposition scheme, an efficient rank search algorithm, low-rank-aware quantization compatibility enhancements, and optimized GPU kernels with operators fusion.
Dong et al. (2024) is an efficient technique for optimizing KV caches in large language models (LLMs). It achieves cache compression by combining sparse strategies with low-rank kernels. Specifically, LESS first employs a sparse strategy to selectively cache a subset of KV pairs based on certain criteria (such as importance or frequency), thereby reducing the size of the cache. Subsequently, it learns the residuals between the original attention outputs and the approximated outputs from the sparse strategy, and accumulates this residual information into a fixed-size low-rank cache. This low-rank cache does not scale with the sequence length, thus effectively reducing memory usage. In this way, LESS not only retains key information but also allows queries to access information discarded by the sparse strategy. As a result, it minimizes performance degradation while maintaining a small cache size. Additionally, LESS can synthesize multiple tokens during generation to produce longer and more coherent sequences. Moreover, both its attention computation and cache update processes are carefully designed.
@@ 4.2.2 KV Merging
    $M(\mathcal{K})=\bigcup_{i=1}^{d}F(\mathcal{S}_{ki})$
    $M(\mathcal{V})=\bigcup_{i=1}^{d}F(\mathcal{S}_{vi})$
@@ 4.2.3 Multi-Level KV Compression
He et al. (2024) propose the Hierarchical Memory Transformer (HMT), a novel framework that enhances the model’s ability to handle long contexts by emulating human memory behavior. With memory-augmented segment-level recurrence, the authors construct a memory hierarchy by retaining tokens from early input segments, passing memory embeddings along the sequence, and recalling relevant information from the history. Initially, the long input text sequence is divided into multiple segments $X_{i}$, each containing a fixed number of tokens. For the first segment $X_{0}$, the model encodes it using a Transformer encoder to obtain the initial memory:
@@ 4.3.1 Offload
Pan et al. (2024) introduces a novel long-context low-power machine learning inference system called InstInfer. This system improves inference performance while reducing energy consumption and costs by offloading the crucial attention computation tasks during the decoding phase from GPUs to CSDs (Computational Storage Devices), through executing attention computations within in-storage.
@@
4.3.2 Memory Management Integrated with the OS System

### v2

@@ 3.2.4 Handling Contradictory Memories
In long-term conversations, the management and construction of memory is a complex and important task, especially when it comes to conflicting memories. It is commonly believed that avoiding conflicting memories is necessary Kim et al. (2024). However, this view does not entirely conform to the true characteristics of human cognition. Human memory is highly context-dependent. The same person may recall different information in different contexts, and even memories of the same event may vary with changes in context. Therefore, allowing memories with conflicting interpretations to coexist within an individual’s memory system not only aligns with the natural laws of human cognition but also adds richness and authenticity to conversations. To effectively deal with these conflicting memories, the author has proposed several optimization strategies for large-language models to choose from flexibly in long-term conversations. First is the “Memory Resolution” strategy. Inspired by entity resolution methods, this strategy merges conflicting memories into an information-rich sentence based on the context of the memory sources, thereby cleverly resolving these conflicts. Second is the “Memory Disambiguation” strategy. The contradiction between two statements may arise from the absence of context, that is, pragmatic ambiguity. Drawing on entity disambiguation methods, the memory disambiguation strategy clarifies the true meaning of each memory by extracting relevant information from the context of the memory, thereby eliminating ambiguity. Finally, there is the “Retention” strategy. Due to the limitations of natural language inference (NLI) models, sometimes memories predicted to be contradictory may not actually be so, but rather a deviation in the model’s understanding. In such cases, retaining these seemingly conflicting memories can provide more possibilities and depth to the conversation. Through these strategies, the author aims to transform conflicting memories into sentences containing richer information about the speaker, thereby enhancing the quality and coherence of long-term conversations.
@@ 3.3 Memory Utilization
The overall structure of Text-based memory utilization is shown in Figure 5. The classification in the figure is based on the search method. Each method is introduced below.
@@ 3.3.1 Full-text Search
Full-text search Chen et al. (2024b) involves scanning the entire text dataset to locate specific character sequences or strings. This method not only looks for exact matches but also considers approximate matches based on text structure and wording. In MAS, when facing broad queries or when the exact location or label of information is unknown, the agent can use a full-text search. For example, when a user asks a broad question about “the impact of climate change,” the agent can utilize a full-text search to browse its entire memory and extract relevant paragraphs or interactions.
@@ 3.3.5 Hash-based Search
Liu et al. (2023) proposed hash-based search methods. In the retrieval task based on memory storage, in order to obtain the most relevant thoughts, the author has designed a two-stage retrieval paradigm. When faced with a new query $Q$, the first step is to obtain the embedding vector $x$ of the query using a large language model agent. This model is capable of mapping the query text into a high-dimensional vector space to capture its semantic information, providing a foundation for the subsequent retrieval process. Subsequently, the retrieval process based on Locality-Sensitive Hashing (LSH) commences. LSH is an efficient approximate nearest neighbor search technique that uses specific hash functions to map similar data points into nearby hash buckets, thereby reducing the number of data points that need to be compared. The design of the hash functions is the core of LSH, and different methods may employ different function designs, such as those based on random projection. Typically, multiple hash functions are used to enhance the accuracy and robustness of the retrieval.
During the retrieval, the hash value of the query vector $x$ is calculated, and the hash buckets that are the same or similar to it are searched for. The data points in these buckets are the candidate’s similar data points. This method can quickly locate the content semantically similar to the query without the need to compare the entire dataset one by one. However, LSH provides approximate results, which may have some errors. But by reasonably designing the hash functions and adjusting parameters, such as the number of hash tables and the combination of hash functions, the errors can be reduced to a certain extent while maintaining the efficiency of the retrieval. The LSH function is as follows,
$\text{hash}(x)=\text{LSH}(x)$ (4)
@@ 3.3.6 Multiple Searches
In the process of retrieval of memories, conducting multiple searches is indeed a good strategy. Through repeated searches, the search scope can be gradually narrowed down, and the accuracy of retrieval can be improved, thereby better meeting the user’s needs. For example, in some complex query scenarios, the initial search may only find part of the relevant information. However, by conducting multiple searches and combining the results of the previous search, the query conditions can be further optimized to obtain more comprehensive and accurate information. In addition, multiple searches can also help language models better understand the user’s needs, thereby generating answers that are more in line with user expectations.
Praas (2023) employs a dense retrieval method, using the SimCSE model to encode both user input and each memory record into vectors. By calculating the cosine similarity, the K-relevant top records are selected for response generation. However, since the retrieved memory records may not all be useful and could even be completely irrelevant, the author introduces a self-reflection mechanism for memory. Specifically, after obtaining the retrieval results, the language model is first required to determine whether this information is sufficient to respond to the user’s input. If it is sufficient, the language model extracts the relevant parts from the memory records as output. If not, the original query needs to be improved by generating keywords or phrases that represent the missing information, which are then expanded into the original query for another round of retrieval. This self-reflection process can be repeated multiple times until the language model finds sufficient information.
@@ 4 KV Cache-Based Memory
This chapter introduces long-term memory based on the KV cache. Since transformer-based LLMs are currently the mainstream, especially with the dominance of the decoder-only structure, KV cache-based memory has become a hot research topic, particularly in the field of long context. In terms of the structure of this chapter, we follow the structure of Chapter 3, introducing memory from three aspects: acquisition, management, and utilization. Since there are many research works related to memory acquisition, we divide it into two sections: KV selection and KV compression. KV selection refers to choosing some KVs to save and discarding the rest. We categorize the selection strategies into eight types: regularity-based summarization, score-based methods, special token embeddings, learning-based KV selection, different selection strategies for different layers and heads, identical KV cache for different layers and heads, locality-sensitive hashing strategy, and backtracking-based KV selection. For KV compression in memory acquisition, we summarize it as low-rank compression, KV merging, multimodal compression, KV quantization, and multi-level compression.
For the second major part, memory management, we divide it into offload strategies (that is, placing memory on cheaper memory), optimization strategies combined with the OS system, storage data structures, and shared attention. We also introduce an article that explores Hash-based Search.
@@
As shown in Figure 6, KV acquisition can be divided into two parts: KV selection and KV compression. KV selection means selecting the most valuable token and evicting other tokens. KV selection means selecting the most valuable token and expelling other tokens.
@@ 4.1.1 Strategy of Regularity
Liu et al. (2024b) speculate that not all tokens need to be stored in memory for large language models to understand context. Just as humans can skim an article and grasp its main ideas, large language models may also be able to skim and understand. It is commonly observed that the attention scores of a token follow a strong power-law distribution, meaning that a token only pays high attention to a few other tokens. More importantly, the authors have observed repeated attention patterns among different tokens in a sequence in trained large language models. Some tokens are more important throughout a paragraph. Specifically, for two different tokens, there is a similarity in the objects they highly attend to, as well as in the objects they ignore. Inspired by these observations, the authors propose the hypothesis of importance persistence: Only those tokens that have had a significant impact in the previous step will have a notable impact in future steps. If this hypothesis holds true, it implies that it is possible to anticipate which tokens may be important for future generations.
@@ 4.1.3 Score Strategy
One important feature of the Quest algorithm Tang et al. (2024b) is its ability to dynamically adjust the contents of the KV cache. Specifically, as the query vector Q changes, Quest re-estimates the criticality of tokens and updates the contents of the KV cache accordingly. This dynamic adjustment ensures that the KV cache always contains the key information required for the current query. The authors’ insight is that to avoid missing critical tokens, they should select pages that contain tokens with the highest attention weights. However, to efficiently select pages, they should compute an approximate attention score based on this insight. The authors found that the upper bound of attention weights within a page can be used to approximate the highest attention within that page. The upper bound of attention weights can be calculated using the channel-wise minimum (mi) and maximum (Mi) of the key vectors. Given a query vector Q, Quest computes the maximum possible value for channel i by taking Ui = max(Qimi, QiMi). Note that Ui is always greater than the product of Qi and any key value Ki on that page, regardless of the sign of Qi. Therefore, when the authors sum Ui, they obtain the upper bound of attention weights for all key vectors on that page. After deriving the upper bound of attention weights, the authors select the top K pages as key pages, where K is an arbitrarily defined hyperparameter. To demonstrate the feasibility of Quest, the authors performed actual self-attention calculations and collected the Top-K attention scores for each page. As shown in Figure 3, the authors’ query-aware sparsity mostly aligns with the predicted sparsity. Quest only performs normal self-attention on the selected pages, which significantly reduces memory movement. The authors define the number of tokens in the selected pages as the "token budget."
@@
Sparse Window Attention (SWA) Zhao et al. (2024) is a method used in the inference process of large language models (LLMs), aiming to reduce memory usage by generating sparse attention patterns while maintaining model accuracy. The core of SWA lies in combining local static and global dynamic sparse patterns: it generates static patterns on local tokens, retaining the most recent tokens to maintain the sequential semantics of the language while capturing the dynamically changing semantic importance of previous tokens through dynamic patterns. This hybrid pattern can more effectively capture key tokens in the sequence.
@@
4.1.4 Special Token Embedding Strategy
@@
4.1.5 Learning-based Token Selection Strategy
@@
Wang et al. (2024b) mainly introduces an algorithm called SQUEEZEATTENTION, which is used to manage KV caches in LLM (Large Language Model) inference to further reduce the memory consumption of inference. The paper points out that existing KV cache optimization methods mostly focus on optimizing from the sequence dimension or the batch dimension while neglecting the potential opportunities in the attention layer dimension. By analyzing the behavior of different attention layers during inference, the authors propose a method for precise re-allocation of cache budgets for attention layers, thereby further reducing the total amount of KV caches.
@@
Fu et al. (2024) propose a Mixture of Attention (MoA), which automatically customizes different sparse attention configurations for various heads and layers. MoA constructs and navigates a search space of various attention patterns and their scaling rules relative to input sequence lengths. It profiles the model, evaluates potential configurations, and identifies the optimal sparse attention compression plan. MoA adapts to different input sizes, revealing that some attention heads expand their focus to accommodate longer sequences, while other heads consistently concentrate on fixed-length local contexts.
@@
SnapKV Li et al. (2024) is a KV cache compression method for optimizing large language models (LLMs) when processing long contexts. It is implemented through the following steps: First, an observation window located at the end of the prompt is used to monitor the model’s attention patterns on input tokens during the generation process, which are stable and consistent throughout the generation. Next, a voting algorithm based on the observation window is employed to identify important positions that are consistently attended to by attention heads during generation, thereby selecting the KV pairs corresponding to the most important tokens. Before voting, SnapKV introduces a pooling-based clustering technique to capture complete local information and filter out irrelevant tokens, thus compressing the context while preserving its integrity. Finally, the selected KV pairs are concatenated with the observation window to form a new, smaller KV cache. The model only needs to perform attention calculations on this significantly reduced cache during subsequent generations, thereby improving efficiency.
@@
4.1.7 Different layers and different heads use the same token selection strategy
@@ 4.2.1 Low-Rank Compression
KV-Cache compression methods typically sample a subset of effective tokens or quantize the data into lower numerical bit widths. However, these methods cannot exploit the redundancy in the hidden dimension of the KV tensors. Chang et al. (2024) presents a hidden dimension compression approach called Palu, a novel KV-Cache compression framework that leverages low-rank projection. Palu decomposes linear layers into low-rank matrices, caches the compressed intermediate states, and dynamically reconstructs the full keys and values. To improve accuracy, compression rate, and efficiency, Palu further includes a medium-grained low-rank decomposition scheme, an efficient rank search algorithm, low-rank-aware quantization compatibility enhancements, and optimized GPU kernels with operator fusion.
Dong et al. (2024) is an efficient technique for optimizing KV caches in large language models (LLMs). It achieves cache compression by combining sparse strategies with low-rank kernels. Specifically, LESS first employs a sparse strategy to selectively cache a subset of KV pairs based on certain criteria (such as importance or frequency), thereby reducing the size of the cache. Subsequently, it learns the residuals between the original attention outputs and the approximated outputs from the sparse strategy and accumulates this residual information into a fixed-size low-rank cache. This low-rank cache does not scale with the sequence length, thus effectively reducing memory usage. In this way, LESS not only retains key information but also allows queries to access information discarded by the sparse strategy. As a result, it minimizes performance degradation while maintaining a small cache size. Additionally, LESS can synthesize multiple tokens during generation to produce longer and more coherent sequences. Moreover, both its attention computation and cache update processes are carefully designed.
@@ 4.2.2 KV Merging
    $M(\mathcal{K})=\bigcup_{i=1}^{d}F(\mathcal{S}_{ki})$ (12)
    $M(\mathcal{V})=\bigcup_{i=1}^{d}F(\mathcal{S}_{vi})$ (13)
@@ 4.2.3 Multi-Level KV Compression
He et al. (2024) propose the Hierarchical Memory Transformer (HMT), a novel framework that enhances the model’s ability to handle long contexts by emulating human memory behavior. With memory-augmented segment-level recurrence, the authors construct a memory hierarchy by retaining tokens from early input segments, passing memory embeddings along the sequence, and recalling relevant information from history. Initially, the long input text sequence is divided into multiple segments $X_{i}$, each containing a fixed number of tokens. For the first segment $X_{0}$, the model encodes it using a Transformer encoder to obtain the initial memory:
@@ 4.3 Memory Management
KV-cache-based memory management can be divided into four parts: offload, integrated with the OS system, shared attention, and the relation between KV cache compression and attention, as shown in Figure 7.
@@ 4.3.1 Offload
Pan et al. (2024) introduces a novel long-context low-power machine learning inference system called InstInfer. This system improves inference performance while reducing energy consumption and costs by offloading the crucial attention computation tasks during the decoding phase from GPUs to CSDs (Computational Storage Devices), by executing attention computations within in-storage.
@@
4.3.2 Memory Management with OS System